# Language Models use Lookbacks to Track Beliefs

### v1

@@ Appendix B The CausalToM Dataset
In total, there are 4 templates (one without and 3 with explicit visibility statements). Each template allows 4 different types of questions (CharacterX asked about ObjectY). We used lists of 103 characters, 21 objects, and 23 states. In our interchange intervention experiments, we randomly sample 80 pairs of original and counterfactual stories.
@@ Appendix F Desiderata-based Component Masking
While interchange interventions on residual vectors reveal where a causal variable might be encoded in the LM’s internal activations, they do not localize the variable to specific subspaces. To address this, we apply the Desiderata-based Component Masking technique (De Cao et al., 2020; Davies et al., 2023; Prakash et al., 2024), which learns a sparse binary mask over the singular vectors of the LM’s internal activations. First, we cache the internal activations from $500$ samples at the token positions where residual-level interchange interventions align with the expected output. Next, we apply Singular Value Decomposition to compute the singular vectors, which are then used to construct a projection matrix. Rather than replacing the entire residual vector with that from the counterfactual run, we perform subspace-level interchange interventions using the following equations:
$\mathsf{W_{proj}=V*V^{T}}$ (2)
$\mathsf{h_{org}\leftarrow W_{proj}h_{counterfactual}+(I-W_{proj})h_{original}}$ (3)
Here, $V$ is a matrix containing stacked singular vectors, while $h_{counterfactual}$ and $h_{original}$ represent the residual stream vectors from the counterfactual and original runs, respectively. The core idea is to first remove the existing information from the subspace defined by the projection matrix and then insert the counterfactual information into that same subspace using the same projection matrix. However, in DCM, instead of utilizing the entire internal activation space, we learn a binary mask over the matrix containing stacked singular vectors to identify the desired subspace. Specifically, before computing the projection matrix, we use the following equations to select the relevant singular vectors:
$\mathsf{V\leftarrow V*mask}$ (4)
We train the mask on $80$ examples of the same counterfactual dataset and use another $80$ as the validation set. We use the following objective function, which maximizes the logit of the causal model output token:
$\mathcal{L}=-\mathsf{logit_{expected\_output}}+\lambda\sum\mathsf{W}$ (5)
Where $\lambda$ is a hyperparameter used to control the rank of the subspace and $\mathsf{W}$ is the parameter of the learnable mask. We trained it for one epoch with ADAM optimizer, a batch size of $4$ and a learning rate of $0.01$.
@@ Appendix G Aligning Character and Object OIs
As mentioned in section 4, the source information, consisting of character and object OI, is duplicated to form the address and pointer of the binding lookback. Here, we describe another experiment to verify that the source information is copied to both the address and the pointer. More specifically, we conduct the same interchange intervention experiment as described in Fig. 5, but without freezing the residual vectors at the state tokens. Based our hypothesis, this intervention will not be able to change the state of the original run, since the intervention at the source information will affect both address and pointer, hence making the model form the original QK-circuit.
@@
Figure 12: Source Information of Binding lookback: In this interchange intervention experiment, the source information—i.e., the character and object OIDs (Refer to caption, Refer to caption)—is modified, while the address and payload (Refer to caption, Refer to caption, Refer to caption) are recomputed based on the modified source. Since both the address and pointer information are derived from the altered source, the binding lookback ultimately retrieves the same original state token as the payload. As a result, we do not observe high intervention accuracy.
@@ Appendix I Speculated Payload in Visibility Lookback
This section presents preliminary evidence that the mechanisms outlined in Sections 4 and 5 generalize to other benchmark datasets. Specifically, we demonstrate that Llama-3-70B-Instruct answers the belief questions (true belief and false belief) in the BigToM dataset Gandhi et al. (2024) in a manner similar to that observed for CausalToM: by first converting token values to their corresponding OIs and then performing logical operations on them using lookbacks. However, as noted in Section 3, BigToM—like other benchmarks—lacks the coherent structure necessary for causal analysis. As a result, we were unable to replicate all experiments conducted on CausalToM. Thus, the results reported here provide only preliminary evidence of a similar underlying mechanism.
@@
Figure 26: Source Information of Binding lookback without freezing address and payload

### v2

@@ Appendix B The CausalToM Dataset
We needed to construct a new dataset because we required a task that models could reliably solve. In contrast, most existing ToM datasets remain challenging for LMs. Additionally, we needed a dataset in which each sample is paired with multiple counterfactuals, enabling causal computations and the extraction of the underlying mechanism. The only dataset that met both criteria was BigToM, which we used in our study. However, even BigToM was insufficient for investigating the full range of factors influencing the mechanism, such as the relationship between a character and their object. Hence, we needed to simplify the task to allow for additional counterfactuals. To test the effect of a specific element, we required the ability to modify only that element without altering the rest of the story or creating an incoherent scenario. For example, consider a BigToM story where a flood occurs, and opening a gate releases the water. In the counterfactual scenario where the gate remains closed, the story’s continuation becomes unintelligible, with the occurrence of a flood.
To address this, we developed CausalToM, which features simple stories accompanied by a range of counterfactuals. Key features include: (1) two characters, objects, and states, (2) the ability to modify each of them independently, and (3) control over whether characters witness each other’s actions. The dataset comprises four templates, one without visibility statements and three with explicit visibility statements. Each template supports four types of questions (e.g., “CharacterX asked about ObjectY”). We used lists of 103 characters, 21 objects, and 23 states. For our interchange intervention experiments, we randomly sampled 80 pairs of original and counterfactual stories.
@@ Appendix D Desiderate Based Patching Via Causal Abstraction
Aligning Features to Causal Variables    In our experiments, we use Singular Vector Decomposition (SVD) to featurize residual stream vectors, i.e., features are the orthogonal singular vectors. For a given transformer layer and token location, we collect the residual stream vectors across a large number of examples and compute the singular vectors. Given singular vector features $\mathbf{F}_{\mathbf{h}}$ of a hidden vector $\mathbf{h}$ in the residual stream of the LM $\mathcal{L}$, we select features to align with a causal variable $A$ in causal model $\mathcal{H}$ using Desiderata-based Component Masking (DCM) (De Cao et al., 2020; Davies et al., 2023; Prakash et al., 2024). Given original input $\mathbf{o}$ and counterfactual input $\mathbf{c}$, we train a mask $\mathbf{m}\in[0,1]^{|\mathbf{F}_{\mathbf{h}}|}$ on the following objective
$\mathsf{CE}\Big(\mathsf{GetLogits}\big(\mathcal{L}_{\mathbf{F}_{\mathbf{h}}\leftarrow\mathbf{m}\circ\mathsf{Get}(\mathcal{L}(\mathbf{c}),\mathbf{F}_{\mathbf{h}}))}(\mathbf{b})\big),\mathsf{GetLogits}\big(\mathcal{H}_{A\leftarrow\mathsf{Get}(\mathcal{H}(\mathbf{c}),A))}(\mathbf{b})\big)\Big)$ (2)
@@ Appendix F Desiderata-based Component Masking
While interchange interventions on residual vectors reveal where a causal variable might be encoded in the LM’s internal activations, they do not localize the variable to specific subspaces. To address this, we apply the Desiderata-based Component Masking technique (De Cao et al., 2020; Davies et al., 2023; Prakash et al., 2024), which learns a sparse binary mask $\mathbf{m}$ over the singular vectors of the LM’s internal activations. We first cache the internal activations from $500$ samples at the token positions specified in the main text for each experiment. Next, we apply Singular Value Decomposition to compute the singular vectors as a matrix $V\in\mathbb{R}^{d\times 500}$ where $d$ is the dimensionality of the residual stream. We then masked this matrix using a learnable binary vector $\mathbf{m}\in[0,1]^{d}$ to choose a subset of singular vectors
$V_{masked}=V\mathbf{m}$ (3)
The chosen subset of vectors is used to construct a projection matrix $W_{proj}\in\mathbb{R}^{d\times d}$.
$W_{proj}=V_{masked}V_{masked}^{T}$ (4)
Then, we perform subspace-level interchange interventions (rather than replacing the entire residual vector) using the following equations:
$h_{new}=W_{proj}h_{c}+(I-W_{proj})h_{o}$ (5)
where $h_{o}$ is the full residual stream of the original run, $h_{c}$ is the full residual stream of the counterfactual run, and $h_{new}$ is the intervened vector where the chosen subspace of $h_{o}$ is replaced with that of $h_{c}$.
The core idea is to first remove the existing information from the subspace defined by the projection matrix and then insert the counterfactual information into that same subspace using the same projection matrix.
In order to find the optimal subspace, we optimize $\mathbf{m}$ to maximize the agreement between the causal model output and the LM’s output. To do so, we train the mask for each experiment on $80$ examples of the same counterfactual datasets specified in the main text and use another $80$ samples as the validation set. We use the following objective function, which maximizes the logit of the causal model output token:
$\mathcal{L}=-\mathsf{logit_{\mathsf{causal\_model\_output\_under\_intervention}}}+\lambda\sum\mathbf{m}$ (6)
Where $\lambda$ is a hyperparameter used to control the rank of the subspace and $\mathbf{m}$ is the learnable mask. See Appendix D for details on how the causal model output under intervention are computed. We trained $\mathbf{m}$ for one epoch with ADAM optimizer, on batches of size $4$ and a learning rate of $0.01$. During training, the parameters of $\mathbf{m}$ are continuous and constrained to lie within the range $[0,1]$. To enforce this constraint, we clamp their values after each gradient update. During evaluation, we binarize the mask by rounding each parameter to the nearest integer, i.e., 0 or $1$.
@@ Appendix G Aligning Character and Object OIs
As mentioned in section 5, the source reference information, consisting of character and object OI, is duplicated to form the address and pointer of the binding lookback. Here, we describe another experiment to verify that the source information is copied to both the address and the pointer. More specifically, we conduct the same interchange intervention experiment as described in Fig. 6, but without freezing the residual vectors at the state tokens. Based on our hypothesis, this intervention will not be able to change the state of the original run, since the intervention at the source information will affect both address and pointer, hence making the model form the original QK-circuit.
@@
Figure 13: Source Reference Information of Binding lookback: In this interchange intervention experiment, the source information, i.e., the character and object OIDs (Refer to caption, Refer to caption), is modified, while the address and payload (Refer to caption, Refer to caption, Refer to caption) are recomputed based on the modified source. Since both the address and pointer information are derived from the altered source, the binding lookback ultimately retrieves the same original state token as the payload. As a result, we do not observe high intervention accuracy.
@@ Appendix I Speculated Payload in Visibility Lookback
This section presents preliminary evidence that the mechanisms outlined in Sections 5 and 6 generalize to other benchmark datasets. Specifically, we demonstrate that Llama-3-70B-Instruct answers the belief questions (true belief and false belief) in the BigToM dataset Gandhi et al. (2024) in a manner similar to that observed for CausalToM: by first converting token values to their corresponding OIs and then performing logical operations on them using lookbacks. However, as noted in Section 3, BigToM—like other benchmarks—lacks the coherent structure necessary for causal analysis. As a result, we were unable to replicate all experiments conducted on CausalToM. Thus, the results reported here provide only preliminary evidence of a similar underlying mechanism.
@@
Figure 27: Source Reference Information of Binding lookback without freezing address and payload